Can someone show me how to find the indicated outputs for f(x)= 4xsquared - 2x?
f(0)=
f(-1) =
f(2) =

Answer by Alan3354(69443)

Can someone show me how to find the indicated outputs for f(x)= 4x^2 - 2x?
f(0)=
Sub 0 for x
f(0) = 4*0^2 - 2*0 = 0
---------------
f(-1) =
Sub -1 for x
f(-1) = 4*(-1)^2 - 2(-1) = 4+2 = 6
-------------
f(2) =
Sub 2 for x